OBSERVATION OF A PLANAR ISOMER OF THE N_{2}O-(C_{2}H_{2})_{2} TRIMER

\underline{S. SHEYBANI-DELOUI}, J. NOROOZ OLIAEE, M. REZAEI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Drive North West, Calgary, Alberta, Canada T2N 1N4; A.R.W. McKELLAR, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

Infrared spectra of a planar isomer of the N_{2}O-(C_{2}H_{2})_{2} are observed in the region of ν_{1} fundamental band of N_{2}O monomer (∼2224 cm^{−1}) using a quantum cascade laser to probe a pulsed supersonic slit jet expansion. The band is simulated by an asymmetric top Hamiltonian with rotational constants of A = 2871 MHz, B = 1140 MHz, C = 816 MHz and hybrid a/b-type transitions. It shows a relatively large blue-shift of about 9 cm^{−1} with respect to the N_{2}O monomer band origin. In addition to the normal istopologue, ^{15}N_{2}O-(C_{2}H_{2})_{2} and N_{2}O-(C_{2}D_{2})_{2} are also observed. Here, we present our observation and experimental results which agree fairly well with a semi-empirical calculation which indicates that the lowest energy isomer has a planar structure very similar to that of OCS-(C_{2}H_{2})_{2}.^{a}

aJ. Norooz Oliaee, A.R.W. McKellar, and N. Moazzen-Ahmadi, Chem. Phys. Lett. 512, 167 (2011).